OpenStudy (anonymous):

PLEASE HELP ME UNDERSTAND THIS ! What is the simplified form of the quantity y-squared minus y minus 12 over the quantity y-squared plus 8y plus 15 ?

OpenStudy (anonymous):

(y^2-y-12)/(y^2 +8y + 15) is this how the problem looks?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

What do you get when you factor y^2-y-12

OpenStudy (zzr0ck3r):

factor them both. Hint: they share a common factor of (y+3)

OpenStudy (anonymous):

(y+3) + (y-4) ?

jimthompson5910 (jim_thompson5910):

it should be (y+3)(y-4) you can check by multiplying them out again (y+3)(y-4) = y(y-4) + 3(y-4) (y+3)(y-4) = y^2 - 4y + 3y - 12 (y+3)(y-4) = y^2 - y - 12

jimthompson5910 (jim_thompson5910):

what does y^2 +8y + 15 factor to?

OpenStudy (anonymous):

(y+5)(y+3) ?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

\[\Large \frac{y^2-y-12}{y^2 +8y + 15}\] \[\Large \frac{(y+3)(y-4)}{y^2 +8y + 15}\] \[\Large \frac{(y+3)(y-4)}{(y+3)(y+5)}\] \[\Large \frac{\cancel{(y+3)}(y-4)}{\cancel{(y+3)}(y+5)}\] \[\Large \frac{y-4}{y+5}\]

OpenStudy (anonymous):

oh okay, thank you so much. I really appreciate the help. c;

jimthompson5910 (jim_thompson5910):

yw